Basal energy requirement. The basal energy requirement B is the number of calories that a person needs to maintain the life process. For a 28 yaer-old female with a height of 160 centimeters and a weight of 45 kilograms(kg), B is 1300 calories. If her weight increases to 50kg, then B is 1365 calories. There is a linear equation that expresses B in therms of her weight w. Find the equation and find the basal energy requirement if her weight is 53.2 kg.
Answer by stanbon(75887)

The basal energy requirement B is the number of calories that a person needs to maintain the life process.
For a 28 year-old female with a height of 160 centimeters and a weight of 45 kilograms(kg), B is 1300 calories. Point (45,1300)
------------------------------------------------------
If her weight increases to 50kg, then B is 1365 calories. There is a linear equation that expresses B in terms of her weight w. Point (50,1365)
------------------------------------------------------
Find the equation and find the basal energy requirement if her weight is 53.2 kg.
---------------------
slope = (1365-1300)/(50-45) = 65/5 = 13
----
intercept: 1300 = 13*45 + b
b = 715
-------------
Equation:
B = 13wt. + 715
If wt = 53.2, solve for B
B(53.2) = 13*53.2 + 715
B(53.2) = 1406.6
--------------
Cheers,
Stan H.